OpenStudy (anonymous):

On a subway route, station C is located at the midpoint between stations A and D. Station B is located at the midpoint between stations A and C. If the distance between stations A and D is 2.4 kilometers, what is the distance between stations B and D?

OpenStudy (primeralph):

Midpoint of a midpoint = 1/2*1/2 = 1/4. 2.4/4. But that's AB To find BD, subtract that from 2.4 BD = 2.4-2.4/4 = 3/4 *2.4 = 1.8

OpenStudy (anonymous):

Wait what? I'm so confused...

OpenStudy (primeralph):